# Multi-height analysis of rational points of toric varieties

### v1

@@ 2.3.2. The action of the Neron-Severi torus on each coordinate
In this paragraph, we shall describe the action of $T_{\operatorname{NS}}$ on $\mathcal{T}_{\Sigma}$. It is induced by the group morphism $\pi^{*}:T_{\operatorname{NS}}\rightarrow T_{\Sigma(1)}$, dual to the morphism $\pi:\operatorname{Div}_{T}(X)\rightarrow\operatorname{Pic}(X)$.
Recall that we have a group morphism
$\exp:\operatorname{Pic}(X)$ ${}^{\vee}_{\mathbf{R}}\hookrightarrow\mathbf{T}_{\operatorname{NS}}(\mathbf{R})$
@@
To any $T$-invariant divisor $D$, we can associate a character $\chi_{D}:T_{\Sigma(1)}\rightarrow\mathbf{G}_{m}$. The composition
@@
corresponds to the character $\chi_{[D]}$ where $[D]\in\operatorname{Pic}(X)$ is the class of $D\in\operatorname{Div}_{T}(X)$.
For any scheme $S$, for any $\rho\in\Sigma(1)$, the action of $t\in T_{\operatorname{NS}}(S)$ on the coordinate of $\mathbf{A}^{\Sigma(1)}$ corresponding to the edge $\rho$ is given by multiplication by $\chi_{[D_{\rho}]}(t)$ as $\chi_{D_{\rho}}\circ\pi^{*}=\chi_{[D_{\rho}]}$. Then we can deduce the following proposition:
@@
Let $a\in\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$. Then for any $\rho\in\Sigma(1)$, the action of $e^{a}\in T_{\operatorname{NS}}(\mathbf{R})$ on the coordinate of $\mathbf{A}^{\Sigma(1)}$ corresponding to the edge $\rho$ is given by multiplication by $e^{\langle[D_{\rho}],a\rangle}$.
2.3.3. A useful covering of the real variety associated to our toric variety
In this paragraph, we suppose that $X(\mathbf{R})$ is equipped with the real topology. Then we shall show the following theorem:
Theorem 2.18.
@@ 2.3.3. A useful covering of the real variety associated to our toric variety
Let $P\in X(\mathbf{R})$ and let $U$ be an open subscheme of $X_{\mathbf{R}}$ such that $P\in U(\mathbf{R})$ and there exists a section $s:U\rightarrow\mathcal{T}_{\mathbf{R}}$ of $q:\mathcal{T}_{\mathbf{R}}\rightarrow X_{\mathbf{R}}$. Such section exists because by Hilbert’s 90, we have $H^{1}_{\text{et}}(X_{\mathbf{R}},T_{\operatorname{NS}})=H^{1}_{\text{Zar}}(X_{\mathbf{R}},T_{\operatorname{NS}})$. Hence we have a continuous section
$P\in U(\mathbf{R})\longmapsto e^{-h_{\mathcal{T},\infty}(s(P))}.s(P)$
[…]
@@
$(P,t)\longmapsto t.e^{-h_{\mathcal{T},\infty}(s(P))}.s(P).$
@@
By [5, Chapitre I, §4, n°5, théorème 1], the map $h_{\mathcal{T},\infty}^{-1}(\{0\})\rightarrow X(\mathbf{R})$ is proper. Moreover, $X(\mathbf{R})$ is compact for the real topology as it is a projective variety. Hence we get our corollary. ∎
2.4. Measure over the universal torsors and Tamagawa number of our variety
@@ 2.4.1. Adelic measure and Tamagawa number over a variety
Here $V$ means a smooth, separated scheme which have geometrically connected fibers over $\operatorname{Spec}(\mathbf{Z})$ and such that $V_{\mathbf{Q}}$ is a smooth, projective and geometrically integral variety over $\mathbf{Q}$. Let us recall here how an adelic metric $(||-||_{v})_{v\in M_{\mathbf{Q}}}$ over $\omega_{V}^{-1}$ defines a measure $\mu_{V}$ over
@@
As in [13, 2.2.1], we can define locally a measure over $V(\mathbf{Q}_{v})$ by the relation
$\|\frac{\partial}{\partial x_{1,v}}\wedge..\wedge\frac{\partial}{\partial x_{n,v}}\|_{v}dx_{1,v}..dx_{n,v}$
@@
By [13, lemma 2.1.2], we have that if $V$ has good reduction at every finite places and the norm is defined by the model (for example if it is a split toric variety), then:
@@
In order to define the Tamagawa number of our variety $V$, one adds some convergence factors $\lambda_{v}=L_{v}(1,\operatorname{Pic}(\overline{V}))$ for $v$ a finite place and $\lambda_{v}=1$ if $v$ is infinite.
@@
The Tamagawa number of $V$ is here defined as $\tau(V)=\mu_{V}(V(A_{\mathbf{Q}}))$ where $\mu_{V}=\prod\limits_{v\in M_{\mathbf{Q}}}\lambda_{v}^{-1}\mu_{V,v}$.
@@
$\lambda_{v}=(1-\frac{1}{p})^{t}.$
@@
$\tau(V)=\mu_{V,\infty}(V(\mathbf{R}))\prod\limits_{p\in\mathbb{P}}(1-\frac{1}{p})^{t}\frac{\sharp V(\mathbf{F}_{p})}{p^{\dim(V)}}.$
Before defining a natural measure over universal torsor, let us recall as in [14, definition 3.1.2], that one can define a natural measure over a split torus $T$:
@@ 2.4.2. Adelic measure over universal torsors
In order to make the notations less cluttered, for the rest of this article, we shall write $\mathcal{T}=\mathcal{T}_{\Sigma}$ where $\mathcal{T}_{\Sigma}$ is the universal torsor defined in 2.11 over the toric variety $X$. In this paragraph, we recall the construction from [17, 4.3.2] to define for any $v\in M_{\mathbf{Q}}$, a measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$.
@@
$1\rightarrow\mathbf{G}_{m,X}\rightarrow q_{*}G_{m,\mathcal{T}}\rightarrow\widehat{T_{\operatorname{NS}}}\rightarrow 0$
associated to the universal torsor, which corresponds up to sign to the element $[\mathcal{T}]\in H^{1}(X,T_{\operatorname{NS}})$ seen as an element in $\operatorname{Ext}^{1}_{X}(\mathbf{G}_{m,X},T_{\operatorname{NS}})$ (see [6, 1.4.3]), we get that $H^{0}(\mathcal{T},\mathbf{G}_{m})=\mathbf{Q}^{\times}$ and that $\operatorname{Pic}(\mathcal{T})$ is trivial. Moreover, by [15, lemme 2.1.10], $\omega_{\mathcal{T}}$ is isomorphic to the pull-back of $\omega_{X}$. Hence there exists a global section of $\omega_{\mathcal{T}}$ corresponding to a trivialization of the line bundle so we get the following proposition [17, proposition 4.24]:
@@
We follow [17, construction 4.25] in order to describe locally the measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$. Let us choose a non zero section $\omega$ of $\omega_{\mathcal{T}}$. For any place $v\in M_{\mathbf{Q}}$, the expression
@@
In our case, recall that $X$ is a smooth, projective and split toric variety and $\mathcal{T}$ an open subset of $\mathbf{A}^{\Sigma(1)}$, the previous construction gives us that $\omega_{\mathcal{T},v}$ is the restriction of the normalised Haar measure $dx_{v}$ over $\mathbf{Q}_{v}^{\Sigma(1)}$ to the open subset $\mathcal{T}(\mathbf{Q}_{v})$ if $v\in M_{\mathbf{Q}}^{0}$ and is the restriction of the natural Lesbesgue measure $dx_{\infty}$ over $\mathbf{R}^{\Sigma(1)}$ to $\mathcal{T}(\mathbf{R})$.
Then using the basic property of the universal torsor, we have the following theorem:
@@
$\omega_{\mathcal{T},v}(\mathcal{T}(\mathbf{Z}_{v}))=\frac{\sharp\mathcal{T}(\mathbf{F}_{v})}{p_{v}^{\dim(\mathcal{T})}}=(1-p_{v})^{t}.\frac{\sharp X(\mathbf{F}_{v})}{p_{v}^{\dim(X)}}.$
@@
for a compact subset $\mathrm{D}\subset\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$. We shall connect it to $\nu(\mathrm{D})$ where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 1.8. We shall need the following lemma which was stated in [14, 4.4.1]:
@@
$P\longmapsto\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt)$
@@
$\int\limits_{\mathcal{T}(\mathbf{R})}f(y)\omega_{\mathcal{T},\infty}(dy)=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt).$
@@
The first assertion is a consequence of the invariance of the Haar measure (see [4, Chapitre VII, §2, n°2, proposition 1]). To show the second assertion, one must remark that in [18], Salberger shows that if you define the measure
@@
it is a $T_{\operatorname{NS}}(\mathbf{R})$-invariant measure and the quotient measure $\frac{\omega_{\text{Sal},\infty}}{\omega_{T_{\operatorname{NS}}}}$ is $\mu_{X,\infty}$. Hence by [4, Chapitre VII, §2, n°2, proposition 4], we have that for any $f:X(\mathbf{R})\rightarrow\mathbf{R}$ compactly supported function, we may choose a section $P\mapsto y_{P}$ such that
@@
$=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}H_{\mathcal{T},\infty}(t.y_{P},\omega_{X}^{-1})f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt)$
$=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt).$
@@
$\omega_{\mathcal{T},\infty}(\mathcal{D})=\int\limits_{P\in X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}\mathbf{1}_{\log_{T_{\operatorname{NS}},\infty}^{-1}(\mathrm{D})}(t)\omega_{T_{\operatorname{NS}}}(t)$
@@
[…]

### v2

@@ 3.3.2. The action of the Neron-Severi torus on each coordinate
In this paragraph, we shall describe the action of $T_{\operatorname{NS}}$ on $\mathcal{T}_{\Sigma}$ seen as the universal torsor (see proposition 3.13). It is induced by the opposite action of the one given by the group morphism $\pi^{*}:T_{\operatorname{NS}}\rightarrow T_{\Sigma(1)}$, dual to the morphism $\pi:\operatorname{Div}_{T}(X)\rightarrow\operatorname{Pic}(X)$ which maps $D$ to $[D]$.
We shall use the exponential group morphism:
$e^{*}:\operatorname{Pic}(X)$ $^{\vee}_{\mathbf{R}}\hookrightarrow\mathbf{T}_{\operatorname{NS}}(\mathbf{R})$
@@
Proposition 3.19.
Let $a\in\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$. For any $\rho\in\Sigma(1)$, the action of $e^{a}\in T_{\operatorname{NS}}(\mathbf{R})$ on the coordinate of $\mathcal{T}_{\Sigma}(\mathbf{R})$ corresponding to the edge $\rho$ is given by multiplication by $e^{-\langle[D_{\rho}],a\rangle}$.
Proof.
To any $T$-invariant divisor $D$, we can associate a character $\chi_{D}:T_{\Sigma(1)}\rightarrow\mathbf{G}_{m}$ which maps $y$ to $X^{D}(y)$. The composition
@@
corresponds to the character $\chi_{[D]}$ where $D]=\pi(D)\in\operatorname{Pic}(X)$ is the class of $D\in\operatorname{Div}_{T}(X)$. Hence for any $\rho\in\Sigma(1)$, following proposition 3.13, the action of $t\in T_{\operatorname{NS}}(\mathbf{R})$ on the coordinate of $\mathcal{T}_{\Sigma}(\mathbf{R})$ corresponding to the edge $\rho$ is given by multiplication by $\chi_{[D_{\rho}]}(t)^{-1}$ as $\chi_{D_{\rho}}\circ\pi^{*}=\chi_{[D_{\rho}]}$. ∎
3.3.3. A natural covering of the real variety associated to our toric variety
In this paragraph, $X(\mathbf{R})$ is seen as a real manifold, endowed of its associated topology. Then we shall prove the following result on the local height $h_{\mathcal{T},\infty}$ (see definition 3.6):
@@ 3.3.3. A natural covering of the real variety associated to our toric variety
Let $P\in X(\mathbf{R})$ and let $U$ be an open subscheme of $X_{\mathbf{R}}$ such that $P\in U(\mathbf{R})$ and that there exists a section $s:U\rightarrow\mathcal{T}_{\mathbf{R}}$ of $q:\mathcal{T}_{\mathbf{R}}\rightarrow X_{\mathbf{R}}$. Such section exists because by Hilbert’s 90, we have $H^{1}_{\text{et}}(X_{\mathbf{R}},T_{\operatorname{NS}})=H^{1}_{\text{Zar}}(X_{\mathbf{R}},T_{\operatorname{NS}})$. Hence we have a continuous section
$P\in U(\mathbf{R})\longmapsto e^{h_{\mathcal{T},\infty}(s(P))}.s(P)$
of $h_{\mathcal{T},\infty}^{-1}(\{0\})\xrightarrow{p}X(\mathbf{R})$ because, by proposition 3.8, we have
$H_{\mathcal{T},\infty}(e^{h_{\mathcal{T},\infty}(s(P))}.s(P),L)=e^{-h_{\mathcal{T},\infty}(s(P))(L)}\cdot H_{\mathcal{T},\infty}(s(P),L)=1$
for any $L\in\operatorname{Pic}(X)$. It gives using lemma 3.8 a trivialisation of the covering as follows:
@@
$(P,t)\longmapsto t.e^{h_{\mathcal{T},\infty}(s(P))}.s(P).$
@@
By [6, chapitre I, §4, n°5, théorème 1], the map $h_{\mathcal{T},\infty}^{-1}(\{0\})\rightarrow X(\mathbf{R})$ is proper. Moreover, $X(\mathbf{R})$ is compact for the real topology as it is a projective variety. Hence we get our corollary. ∎
3.4. Measure over the universal torsors and Tamagawa number
@@ 3.4.1. Adelic measure and Tamagawa number over a variety
Let $V$ be a smooth, separated scheme which have geometrically connected fibers over $\operatorname{Spec}(\mathbf{Z})$ and such that $V_{\mathbf{Q}}$ is a smooth, projective and geometrically integral variety over $\mathbf{Q}$. Let us recall here how an adelic metric $(||-||_{v})_{v\in M_{\mathbf{Q}}}$ over $\omega_{V}^{-1}$ defines a measure $\mu_{V}$ over
@@
As in [15, 2.2.1], we can define locally a measure over $V(\mathbf{Q}_{v})$ by the relation
$\Big\|\frac{\partial}{\partial x_{1,v}}\wedge..\wedge\frac{\partial}{\partial x_{n,v}}\Big\|_{v}dx_{1,v}..dx_{n,v}$
@@
By [15, lemma 2.1.2], we have that if $V$ has good reduction at every finite places and the norm is defined by the model (such hypotheses are verified by a split toric variety), then:
@@
In order to define the Tamagawa number of our variety $V,$ we first define convergence factors. We set $\lambda_{v}=L_{v}(1,\operatorname{Pic}(\overline{V}))$ for $v$ a finite place and $\lambda_{v}=1$ for $v$ is infinite.
@@
The Tamagawa number of $V$ is defined as $\tau(V)=\mu_{V}(V(A_{\mathbf{Q}}))$ where $\mu_{V}=\prod\limits_{v\in M_{\mathbf{Q}}}\lambda_{v}^{-1}\mu_{V,v}$.
@@
$\lambda_{v}=\left(1-\frac{1}{p}\right)^{t}.$
@@
$\tau(V)=\mu_{V,\infty}(V(\mathbf{R}))\prod\limits_{p\in\mathbb{P}}\left(1-\frac{1}{p}\right)^{t}\frac{\sharp V(\mathbf{F}_{p})}{p^{\dim(V)}}.$
Before defining a natural measure over universal torsor, let us recall [16, definition 3.1.2], that one can define a natural measure over a split torus $T$:
@@ 3.4.2. Adelic measure over universal torsors
In order to make the notations less cluttered, for the rest of this article, we shall write $\mathcal{T}=\mathcal{T}_{\Sigma}$ where $\mathcal{T}_{\Sigma}$ is the universal torsor defined in proposition 3.13 over the toric variety $X$. In this paragraph, we recall the construction from [19, section 4.3.2] to define for any $v\in M_{\mathbf{Q}}$, a measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$. We write $\widehat{T_{\operatorname{NS}}}$ for the sheaf of characters associated to the Neron-Severi torus.
@@
$1\rightarrow\mathbf{G}_{m,X}\rightarrow q_{*}\mathbf{G}_{m,\mathcal{T}}\rightarrow\widehat{T_{\operatorname{NS}}}\rightarrow 0$
associated to the universal torsor, which corresponds up to sign to the element $\mathcal{T}]\in H^{1}(X,T_{\operatorname{NS}})$ seen as an element in $\operatorname{Ext}^{1}_{X}(\mathbf{G}_{m,X},T_{\operatorname{NS}})$ (see [7, 1.4.3]), we get that $H^{0}(\mathcal{T},\mathbf{G}_{m})=\mathbf{Q}^{\times}$ and that $\operatorname{Pic}(\mathcal{T})$ is trivial. Moreover, by [17, lemme 2.1.10], $\omega_{\mathcal{T}}$ is isomorphic to the pull-back of $\omega_{X}$. Hence there exists a global section of $\omega_{\mathcal{T}}$ corresponding to a trivialization of the line bundle so we get the following proposition [19, proposition 4.24]:
@@
We follow [19, construction 4.25] in order to describe locally the measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$. Let us choose a non zero section $\omega$ of $\omega_{\mathcal{T}}$. For any place $v\in M_{\mathbf{Q}}$, the expression
@@
In our case, recall that $X$ is a smooth, proper and split toric variety and $\mathcal{T}$ an open subset of $\mathbf{A}^{\Sigma(1)}$, the previous construction gives us that $\omega_{\mathcal{T},v}$ is the restriction of the normalised Haar measure $dx_{v}$ over $\mathbf{Q}_{v}^{\Sigma(1)}$ to the open subset $\mathcal{T}(\mathbf{Q}_{v})$ if $v\in M_{\mathbf{Q}}^{0}$ and is the restriction of the natural Lesbesgue measure $dx_{\infty}$ over $\mathbf{R}^{\Sigma(1)}$ to $\mathcal{T}(\mathbf{R})$.
We shall compute $\omega_{\mathcal{T},v}(\mathcal{T}(\mathbf{Z}_{v}))$ for $v\in M_{\mathbf{Q}}^{0}$:
@@
$\omega_{\mathcal{T},v}(\mathcal{T}(\mathbf{Z}_{v}))=\frac{\sharp\mathcal{T}(\mathbf{F}_{v})}{p_{v}^{\dim(\mathcal{T})}}=\left(1-\frac{1}{p_{v}}\right)^{t}.\frac{\sharp X(\mathbf{F}_{v})}{p_{v}^{\dim(X)}}.$
Proof.
Because $\omega_{\mathcal{T},v}$ is the restriction of the normalised Haar measure $dx_{v}$ over $\mathbf{Q}_{v}^{\Sigma(1)}$ to the open subset $\mathcal{T}(\mathbf{Q}_{v})$ , we have $\omega_{\mathcal{T},v}(\mathcal{T}(\mathbf{Z}_{v}))=\frac{\sharp\mathcal{T}(\mathbf{F}_{v})}{p_{v}^{\dim(\mathcal{T})}}$. Using that $T_{\operatorname{NS}}(\mathbf{F}_{v})$ acts freely on $\mathcal{T}(\mathbf{F}_{v})$ and that $X(\mathbf{F}_{v})=\mathcal{T}(\mathbf{F}_{v})/T_{\operatorname{NS}}(\mathbf{F}_{v})$, we can conclude the proof. ∎
@@
for a compact subset $\mathrm{D}\subset\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$. We shall connect it to $\nu(\mathrm{D})$ where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 2.9. We shall need the following lemma which was stated in [16, 4.4.1]:
@@
$P\longmapsto\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}^{-1}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt)$
@@
$\int\limits_{\mathcal{T}(\mathbf{R})}f(y)\omega_{\mathcal{T},\infty}(dy)=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}^{-1}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt).$
@@
The first assertion is a consequence of the invariance of the Haar measure (see [5, chapitre VII, §2, n°2, proposition 1]). To show the second assertion, one must remark that in [20, section 5], Salberger shows that if you define the measure
@@
it is a $T_{\operatorname{NS}}(\mathbf{R})$-invariant measure and the quotient measure $\frac{\omega_{\text{Sal},\infty}}{\omega_{T_{\operatorname{NS}}}}$ is $\mu_{X,\infty}$. Hence by [5, chapitre VII, §2, n°2, proposition 4], we have that for any $f:X(\mathbf{R})\rightarrow\mathbf{R}$ compactly supported function, we may choose a section $P\mapsto y_{P}$ such that
@@
$=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}H_{\mathcal{T},\infty}(t.y_{P},\omega_{X}^{-1})f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt).$
Using that $H_{\mathcal{T},\infty}(t.y_{P},\omega_{X}^{-1})=|\chi_{\omega_{X}^{-1}}(t)|^{-1}_{\infty}.H_{\mathcal{T},\infty}(y_{P},\omega_{X}^{-1})$, we can conclude the proof. ∎
@@
$\omega_{\mathcal{T},\infty}(\mathcal{D})=\int\limits_{P\in X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}^{-1}\mathbf{1}_{\log_{T_{\operatorname{NS}},\infty}^{-1}(\mathrm{D})}(t)\omega_{T_{\operatorname{NS}}}(t)$
@@
$\omega_{\mathcal{T},\infty}(\mathcal{D})$ $=2^{t}.\int\limits_{P\in X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}}e^{\langle\omega_{X}^{-1},a\rangle}\mathbf{1}_{\mathrm{D}}(a)da$